A=Vio . How long will it take $250 to double at 2% compounded quarterly? D ih
Answer #1

We know that A=.P(HOME P250 A = 2P =500 Y = . > 0, 07 n=4 ( quarterly) ta? 500 =250(1+0.02745 2= (1+0.005)45 2= (1.0054t tn (